OpenStudy (anonymous):

Find a quadratic model for the set of values. x: -2 0 4 f(x): 1 -3 85 A. y = 3x^2 + 6x - 4 B. y = -4x^2 - 6x + 3 C. y = 4x^2 + 6x - 3 D. y = 6x^2 + 4x - 3

jimthompson5910 (jim_thompson5910):

since you're given a list of choices, you can plug in each x value and see if you get the correct y value out

jimthompson5910 (jim_thompson5910):

for example, the table states that if x = 0, then f(x) = -3 so plugging x = 0 should give you y = -3 this is the same for each x,y pairing

OpenStudy (calculusxy):

Formula: (x+a)(x+b)=x^2+bx+ax+ab=x^2+(ab)^2+ab